OpenStudy (anonymous):

3^(k+2) can be rewritten into what? 3^k + 3^2 or 3^k * 3^2

jimthompson5910 (jim_thompson5910):

in general a^(b+c) = a^b * a^c

jimthompson5910 (jim_thompson5910):

ex: 2^(3+4) = 2^3*2^4 on the left side you get 2^(3+4) = 2^7 = 128 on the right side you get 2^3*2^4 = 8*16 = 128 so that example supports the general equation above

OpenStudy (anonymous):

yea was able to figure it out by using the number examples thanks

jimthompson5910 (jim_thompson5910):

yw